Five years from now Mary will be twice as old as her daughter Ashley. Currently the sum of their ages is 98. Find their current ages.
Answer by ewatrrr(24785)

Let M be Mary's age and A be Ashley's age Five years from now.
M = 2* A
Question also states the following to is true about now!!
(M-5)+(A-5)= 98
Substitute for M:
2*A-5+ A-5 = 98
Combine like terms and solve:
A=36
M= 2*36 = 72
currently:
Mary is 67 and Ashley is 31
Check Your answer: 67-31 = 98